OpenStudy (anonymous):

Solve the absolute value equation. Check your solutions. | 5x + 13| = –7

OpenStudy (cruffo):

no solution. The absolute value of an expression is always nonnegative.

OpenStudy (anonymous):

can some1 explain how i get the answer

OpenStudy (anonymous):

im lost

OpenStudy (helder_edwin):

as cruffo pointed out. the absolute value of any expression is always positive so a situation like \[ \large |\text{whatever|=-7} \] is impossible to satisfy. so the equation has no solutions.

OpenStudy (anonymous):

omg (5x + 13) = -7 5x = - 7 - 13 5x = - 20 x = - 4 is this right then

OpenStudy (cruffo):

The absolute value of a number is the distance from 0 to the number. For example, |2| = 2 and |-2| = 2

OpenStudy (cruffo):

the symbols | | is not the same as ( )

OpenStudy (anonymous):

okay now i get it thank u guys